### v2

# Iterative Compositional Data Generation for Robot Control

Abstract
Collecting robotic manipulation data is expensive, making it impractical to acquire demonstrations for the combinatorially large space of tasks that arise in multi-object, multi-robot, and multi-environment settings. While recent generative models can synthesize useful data for individual tasks, they do not exploit the compositional structure of robotic domains and struggle to generalize to unseen task combinations. We propose a semantic compositional diffusion transformer that factorizes transitions into robot-, object-, obstacle-, and objective-specific components and learns their interactions through attention. Once trained on a limited subset of tasks, we show that our model can zero-shot generate high-quality transitions from which we can learn control policies for unseen task combinations. Then, we introduce an iterative self-improvement procedure in which synthetic data is validated via offline reinforcement learning and incorporated into subsequent training rounds. Our approach substantially improves zero-shot performance over monolithic and hard-coded compositional baselines, ultimately solving nearly all held-out tasks and demonstrating the emergence of meaningful compositional structure in the learned representations.
Figure 1: Iterative Compositional Data Generation.
1 Introduction
Augmenting model training with self-generated data is a promising approach to improve sample efficiency in domains where collecting real data is expensive. In the context of robotic manipulation, gathering new experience requires operating a physical robot—a process that is labor- and time-intensive, and incurs wear, maintenance, and energy costs. Consequently, collecting data from scratch for every possible new manipulation task quickly becomes impractical as evidenced by various large-scale data collection efforts (walke2023bridgedata; openx2023openx; khazatsky2024droid). Recent work has shown that current generative models can produce data of sufficient quality to enable training models with substantially reduced real-world experience, including in control settings (yu2023scaling; lu2023synthetic; liang2023adaptdiffuser). However, most existing approaches focus on improving sample efficiency within a single task, and do not leverage self-generated data to accelerate learning on entirely new tasks (janner2022planning; lu2023synthetic).
In this work, we investigate whether a robot learning system can iteratively improve its ability to solve unseen tasks by generating artificial training data for those unseen tasks with a self-improving generative model (Figure 1). We leverage the insight that cross-embodiment robotic manipulation domains exhibit an inherent compositional structure, whereby each task solution involves a unique composition of reusable models of objects, skills, and controllers. Our central hypothesis is that constructing model architectures that explicitly exploit this compositionality enables zero-shot generation of high-quality synthetic training data for novel task compositions, mitigating the need to relearn every task from scratch on a physical robot.
We focus on a reinforcement learning (RL) setting in which tasks are defined compositionally (Mendez2022; hussing2024robotic), constructed by combining a small number of elements such as robots, objects, obstacles, and goals. Intuitively, machine learning approaches can exploit the inherent combinatorial structure of these domains to generalize to unseen task configurations. Yet, standard single-and multi-task agents require vast amounts of data in such settings, struggling to exploit the compositional structure when the available data is small. Learners are better able to exploit the structure when the policy architecture mirrors the underlying task factorization (devin2017learning; andreas2017modular; Mendez2022; mendez2022modular).
One outstanding challenge is that pre-defining such architectures requires strong prior knowledge about the correct decomposition. While much prior engineering knowledge is available for the robotics tasks we consider, it is unclear that these priors are optimal. In this work, we instead train a transformer to learn the compositional structure directly from data, leveraging the transformer interpretation as a graph neural network (GNN). Instead of training a policy on a subset of task combinations and evaluating its zero-shot capabilities, we train a diffusion transformer on the same subset of tasks to generate training data for the policies of unseen task combinations, thereby reducing the amount of data required to learn novel behaviors. The model learns a separate tokenizer for each individual task module (e.g., a specific robot, object, or environment) and uses cross-attention to infer the graph that connects these encoders. This yields a representation that is analogous to the hard-coded compositional network used in earlier work (Mendez2022), but the structure is learned from data rather than specified a priori.
We first demonstrate that our task-conditional diffusion transformer enables superior zero-shot generalization capabilities compared to monolithic architectures. We then highlight the need for compositional tokenization by showing that models with factor-specific tokenizers achieve improved zero-shot performance relative to models that rely on a single shared tokenizer. Next, we show that models that properly learn the underlying task decomposition can be iteratively trained on their own generated data for unseen tasks to produce training data for solving more new tasks without requiring additional real data. Finally, we analyze the learned representations and find that the model discovers a decomposition that differs from previous work, indicating that effective compositional structure can emerge automatically from data.
2 Preliminaries
We formulate our problem as generating data of and learning policies in a Markov decision process (MDP) $\mathcal{M}_{n}=(S_{n},A_{n},R_{n},\mathcal{P}_{n},T)$} where $S_{n}$ is the state space, $A_{n}$ is the action space, $R_{n}$ is the reward function mapping state-action pairs $(s,a)$ to a scalar $r$, $\mathcal{P}_{n}$ is the transition probability function determining the next state $s^{\prime}$ given the current state-action pair $(s,a)$, and $T$ is the task horizon. The rewards are bounded to the range $r\in[0,1]$. We say a agent succeeds if, at any state in a trajectory, it achieves the maximum reward of $r=1$. We also define the function $D_{n}:S_{n}\mapsto A_{n}$ which outputs a termination signal $d$ that indicates if the agent has moved to an absorbing, non-rewarding state. We consider a set of $N$ such MDPs $\{\mathcal{M}_{n}\}_{n=0}^{N-1}$ and our goal is to learn a policy $\pi^{*}=\{\pi_{n}^{*}\}_{n=0}^{N-1}$ that maximizes the average probability of success over this set, that is, $\pi^{*}\in\arg\max_{\pi}\frac{1}{N}\sum_{n=0}^{N-1}\mathbb{E}_{\pi_{n},\mathcal{M}_{n}}\big[\mathbf{1}[\max_{0\leq t\leq T}r_{t}=1]\big]$.
2.1 CompoSuite Benchmark
CompoSuite (Mendez2022) is a simulated robotic manipulation benchmark for evaluating compositional reinforcement learning (RL) agents. CompoSuite provides a family of $4\times 4\times 4\times 4=256$ distinct manipulation tasks by composing exactly one out of four elements from each of the following four axes:
  • Robot arms   KUKA’s IIWA, Kinova’s Jaco, Franka’s Panda, Kinova’s Gen3.
  • Objects   Box, Dumbbell, Plate, Hollow box.
  • Obstacles   No obstacle, Wall blocking object, Doorway near object, Wall blocking goal.
  • Objectives   Pick and place, Push, Place on shelf, Place in trash can.
(a) IIWA, Box,
None, PickPlace
(b) Jaco, Dumbbell, ObjectWall, Push
(c) Panda, Plate, ObjectDoor, Shelf
(d) Kinova3, Hollowbox, GoalWall, Trashcan
Figure 2: Four example CompoSuite tasks, each defined by selecting one element from each axis.
Figure 3: Overview of the 16-dimensional task indicator. For every task in CompoSuite, the model receives a binary vector formed by concatenating four one-hot segments: a 4-dimensional robot ID, 4-dimensional object ID, 4-dimensional obstacle ID, and 4-dimensional subtask ID. Each segment activates exactly one entry corresponding to the chosen element along that axis. The example demonstrates how the task Kinova3, Plate, None, Push is encoded into the final 16-element vector.
For each task, states are provided as symbolic representations containing proprioceptive robot features (joint and gripper positions and velocities) together with absolute and relative Cartesian positions of the object, obstacle, and goal in the scene. Each task is defined by selecting exactly one unique element from each of the four axes: Robot, Object, Obstacle, and Objective. To illustrate the compositional structure, Figure 2 shows four example tasks from CompoSuite. The state vector also contains a binary indicator vector of length 16 that identifies a task via four one-hot sub-vectors, one for each axis. Figure 3 illustrates this layout. Rewards are defined with dense, stage-wise rewards to guide the learning.
hussing2024robotic released 1 million transitions for every task in CompoSuite across four dataset variants (approximately 1 billion transitions in total). The four datasets span a range of performance levels, from early training to expert proficiency, and were generated using policies trained with Proximal Policy Optimization (schulman2017ppo) and Soft Actor-Critic (haarnoja2018sac). Trajectories are stored as transition tuples $\langle s,a,r,s^{\prime},d\rangle$. For our experiments, we focus exclusively on the expert datasets.
2.2 Diffusion Models
A diffusion model is a generative model that learns to reverse a gradual noising process applied to data (ho2020denoising). We use diffusion models to generate artificial data for training. More precisely, we use the Elucidated Diffusion framework (Karras2022). Given a data vector $x_{0}\in\mathbb{R}^{d}$, we consider a collection of noise levels $\{\sigma_{t}\}_{t=1}^{T}$ with $\sigma_{t}>0$. For each $t$, the forward process $q$ produces a noised sample $x_{t}$ by adding Gaussian noise of magnitude $\sigma_{t}$:
$x_{t}=x_{0}+\sigma_{t}\varepsilon\kern 5.0pt,\quad\varepsilon\sim\mathcal{N}(0,I)\kern 5.0pt,\quad\log\sigma_{t}\sim\mathcal{N}(P_{\mathrm{mean}},P_{\mathrm{std}}^{2})\kern 5.0pt,\quad t=1,\dots,T\kern 5.0pt,$
so that, for any fixed noise level $\sigma_{t}$, the conditional distribution of $x_{t}$ given $x_{0}$ is $q(x_{t}\mid x_{0},\sigma_{t})=\mathcal{N}\bigl(x_{0},\sigma_{t}^{2}I\bigr).$ Here, $P_{\mathrm{mean}}\in\mathbb{R}$ and $P_{\mathrm{std}}>0$ are scalar hyperparameters that control the mean and standard deviation of the log-noise distribution $\log\sigma_{t}$. A neural network $\varepsilon_{\theta}$ is trained to predict the clean sample $x_{0}$ from a noised sample $x_{t}$ and its associated noise level $\sigma_{t}$. The training objective is a noise-weighted reconstruction loss:
$\mathcal{L}_{\text{diff}}(\theta)=\mathbb{E}_{x_{0}\sim p_{\text{data}},\,\sigma_{t},\,\varepsilon\sim\mathcal{N}(0,I)}\!\left[\left\|\varepsilon_{\theta}(x_{t},\sigma_{t})-x_{0}\right\|_{2}^{2}w(\sigma_{t})\right]\kern 5.0pt,\quad w(\sigma)=\frac{\sigma^{2}+\sigma_{\mathrm{data}}^{2}}{(\sigma\,\sigma_{\mathrm{data}})^{2}}\kern 5.0pt,$
where $\sigma_{\mathrm{data}}>0$ is a hyperparameter representing the typical data scale. At generation time, the model constructs a reverse denoising process over a decreasing sequence of noise levels $\{\sigma_{t}\}_{t=1}^{T}$. Starting from a high-noise initialization $x_{T}\sim\mathcal{N}(0,\sigma_{T}^{2}I)$, the model iteratively applies the denoiser to define reverse transitions $p_{\theta}(x_{t-1}\mid x_{t})$ until it obtains a synthetic sample $x_{0}$ that approximately follows the data distribution.
3 Task-Graph Compositional Transformer for Iterative Data Generation
We assume a functionally compositional task graph. mendez2022modular define a hard-coded set of modules, each representing a task element, and define task solutions as fixed paths through that graph. We instead assume that each task consists of basic elements, each of which is a random variable corresponding to one component of the transition, such as a factor of the state or next state, an action, a reward, or a termination indicator. Let $\mathcal{F}$ denote the set of all such basic elements. Then, each element $f\in\mathcal{F}$ is associated with an input space $X^{f}$ and a representation space $Y^{f}$. An encoder-decoder pair $(e_{f},o_{f})$ maps raw variables into the representation space $e_{f}:X^{f}\mapsto Y^{f}$ and back $o_{f}:Y^{f}\mapsto X^{f}$. We define a computation graph $G=(V,E)$ that captures the shared structure across all tasks, where the vertices $V={Y^{f}}_{f\in\mathcal{F}}$ are the representation spaces and the edges $E$ are transformations that specify how information can flow between representations. A specific MDP $\mathcal{M}_{n}$ is characterized by a subset of elements present in that task, $\mathcal{F}_{n}\subseteq\mathcal{F}$, and the induced subgraph of $G$ on their representation spaces, with a joint distribution over their values. The CompoSuite benchmark fits this view: a task is obtained by selecting one element along the robot, object, obstacle, and objective axes, thereby instantiating a particular set of state-factor vertices and their interactions.
As the graph operates in representation space, it can be used to instantiate a variety of learned functions on an MDP, such as a policy or a conditional generative model. A probabilistic model defined on $G$ can specify a conditional distribution over the unobserved basic elements given the values of any subset of observed ones.
3.1 Transformers as Graphs
Hard-coding the structure of the computation graph requires extensive domain knowledge and may result in a suboptimal architecture. In consequence, we would like to learn the graph structure directly from data. For this, we rely on the finding that the well-known transformer architecture (vaswani2017attention) can be interpreted as a GNN (joshi2025transformers). In particular, a transformer with $L$ layers maps a sequence of input tokens $x_{1},\dots,x_{K}$ to a sequence of output representations $h_{1}^{L},\dots,h_{K}^{L}$ by repeatedly applying self-attention and feed-forward layers. For each token $i$ and layer $\ell$, the model computes queries, keys, and values $q_{i}^{\ell}=W_{Q}h_{i}^{\ell-1}$, $k_{j}^{\ell}=W_{K}h_{j}^{\ell-1}$, $v_{j}^{\ell}=W_{V}h_{j}^{\ell-1}$ where $W_{Q},W_{K},W_{V}$ are learned weight matrices. The model then assigns the $i$-th token’s attention weights over each other token $j$ as $\alpha_{ij}^{\ell}=\mathrm{softmax}j\big(q_{i}^{\ell\top}k_{j}^{\ell}/\sqrt{d}\big)$ and aggregates other tokens’ values into an updated representation $h_{i}^{\ell}=\mathrm{FF}\left(\sum{j=1}^{N}\alpha_{ij}^{\ell}v_{j}^{\ell}\right)$}, where $\mathrm{FF}$ implements a feed-forward layer and $h_{i}^{0}=x_{i}$. In the interpretation of the transformer as a GNN, each token $i$ corresponds to a node with a feature vector $h_{i}^{\ell-1}\in\mathbb{R}^{d}_{h}$, and self-attention implements message passing on a fully connected directed graph over these nodes. By learning the weight matrices $W_{Q},W_{K},W_{V}$, the transformer learns to assign high attention from token $i$ to token $j$ exactly when the graph should contain a strong directed edge from node $j$ to node $i$.
Interpreting the transformer as a GNN suggests that a transformer can automatically discover the underlying graph structure of a set of problems that are related compositionally. This reduces the architectural challenge of designing an appropriate graph to designing a tokenization scheme that enables representing such a graph.
3.2 Semantic Compositional Diffusion Transformers
Thus, we set out to encode our task graph in a diffusion model by implementing $\epsilon_{\theta}$ as a diffusion transformer (DiT; peebles2023scalable). This model processes noised inputs $(x_{t,1},\dots,x_{t,K})$ in the original transition space and outputs denoised predictions $\epsilon_{\theta}(x_{t},t)\in\mathbb{R}^{K\times d_{x}}$ at each diffusion step, which is interpreted as a prediction of the added noise for each component. Our diffusion transformer architecture internally uses factor-specific encoders to map inputs to token embeddings, processes these through self-attention, and decodes back to the original space at each step of the reverse diffusion process.
We construct the input sequence for each transition by associating each component of our task graph $f\in\mathcal{F}$ with the elements of CompoSuite as described in Section 2.1. For both the state and next state, we treat each element per axis as one factor—e.g. each robot arm is one factor. In addition, we add one factor each for the action, reward, and termination signal. This yields a DiT that can learn directly on top of the task graph. For representation learning, we equip each factor with a parametric encoder-decoder pair $(e_{f,\theta},o_{f,\psi})$, both instantiated as neural networks. The encoder maps the inputs into a learned embedding $y^{f}=e_{f,\theta}(z^{f})$, which we interpret as living in the representation space $Y^{f}$. The collection $(y^{f})_{f\in\mathcal{F}}$ is treated as the $K$ tokens processed by the transformer. Self-attention over this factor-specific set of tokens implements graph-compositional inference: at each diffusion step, the representation of every factor $f$ is updated by attending to all other factors $f^{\prime}\in\mathcal{F}$, so that, for example, the current robot token can condition on the current object, obstacle, and objective tokens in a way that mirrors the edges of $G=(V,E)$. At each diffusion step, the transformer outputs denoised token embeddings $\bar{y}^{f}$, which are then mapped back to the original variable domains with the decoders, yielding predictions $(o{f,\psi}(\bar{y}^{f})){f\in\mathcal{F}}$ in the original transition space.
For conditioning, the original DiT injects variables such as the diffusion step $t$ through adaptive layer normalization. This produces per-block scale and shift parameters that gate the self attention and feed-forward updates. We implement our task conditioning via an additional input embedding that modulates all transformer blocks. For each diffusion step $t$ and task index $n$, we form a context embedding $u(t,n)=E_{t}(t)+E_{n}(n)$} and pass it through a small network that produces adaptive normalization and gating parameters for every block of the DiT. This pathway injects $(t,n)$ into the model only through these adaptive transforms, and leaves the compositional semantics of the factor-specific tokenization unchanged. The resulting network provides us with a diffusion model that can be trained to generate RL transitions for each task by simply selecting the correct encoders and conditioning. We visualize our proposed architecture in Figure 4.
Figure 4: Visualization of our semantic compositional transformer architecture. We factorize each transition into state factors, actions, reward, and terminal indicators. State factors represent the compositional dimensions along which tasks can vary (for example, different robots). Each state factor has its own encoder-decoder pair, providing partial parameter separation. The encoded tokens, augmented with positional embeddings, are processed by several diffusion transformer layers. The diffusion transformer uses adaptive statistics conditioned on timestep and task indicator embeddings. Its output tokens are then decoded by the factor-specific decoders, with state encoder-decoder pairs shared between current and next state.
Using this architecture for diffusion modeling induces a joint representation over factor-specific component embeddings. For each task index $n$, the learned diffusion model defines a distribution $p_{\theta}(x_{0}\mid n)$ over denoised transitions $x_{0}$ in the original transition space. Within the denoiser, factor-specific encoders map each component to token embeddings $(\bar{y}^{f}){f\in\mathcal{F}}$, the shared diffusion transformer blocks (modulated by task conditioning through adaptive layer normalization) process these through self-attention to learn relationships between factors, and decoders map the embeddings back to the original space. As this joint representation over component embeddings is shared across tasks, we can use structure learned from one task to improve the marginals for others and incrementally refine each factor’s predictive distribution as new tasks are added.
3.3 Self-Refining Compositional Distributions
We use our trained DiT to produce training data for two purposes: training behavior policies for unseen tasks (for which we do not have real training data), and updating the DiT itself. By virtue of the compositional graph structure, our model can train on a set of compositional tasks, generate data for new task combinations zero-shot, and use the generated data to improve the DiT. Notably, because we learn factorized pieces of a distribution to improve their marginals, the generator improves not only on the tasks for which we generate data, but also on tasks that share some of these factors. As an illustrative example, consider tasks that vary on the robot and object (e.g., a subset of CompoSuite). Suppose the dataset contains transitions for three tasks: (Panda, Box), (Jaco, Box), and (Panda, Plate), but no real data for (Jaco, Plate). Our compositional DiT can generate plausible transitions for (Jaco, Plate) by combining the learned Jaco and Plate factors. Retraining on these additional samples increases the effective data available for the shared Plate factor and constrains it across multiple robotic contexts, so that any downstream task involving Plate —e.g., (IIWA, Plate)—benefits from a sharper object marginal than would be possible from the original tasks alone.
We summarize our procedure in Algorithm 1. The algorithm starts with a set of (real) data from a sub-set of training tasks and proceeds in rounds. In every round, we fit our DiT to all training data. We then generate data for all existing validation and test tasks, train a policy using TD3-BC (Fujimoto2021Minimalist) on the generated data, and evaluate the policy online in the environment. If the success rate (sr) for any task is larger than some threshold $\tau$, we add the data generated for this task to the training set. This performance-based filter only admits synthetic data that is useful for training high-quality policies. If no tasks surpass the quality threshold, we increase a patience parameter $c$. When the patience exceeds a predefined threshold $C$, we decrease the quality threshold $\tau$.
The approach in Algorithm 1 uses self-generated data to train a data generator. One question is whether adding sub-optimal data for a single task might lead to degraded performance across all tasks. Training on iteratively generated data of prior versions of a generative model often results in the performance of the model decreasing over time, a phenomenon known as model collapse (shumailov2024ai). Our compositional transformer architecture from Section 3.2 uses the data generated for a particular task exclusively to train the encoder-decoder pairs specific to that task’s elements. In consequence, each generated dataset contributes only to a subset of all transformer weights (e.g., data generated for the Panda encoder-decoder is not used to update the parameters of the Jaco encoder-decoder). This mechanism inherent to our compositional architecture partially guards our DiT against model collapse.
Algorithm 1 Compositional Iterative Bootstrapping with Synthetic Data
1:Initial dataset $\mathcal{D}$; target task set $\mathcal{T}_{\text{target}}$; initial threshold $\tau_{0}$ with lower bound $\tau_{\min}$; threshold step size $\Delta_{\tau}$; patience $C$; maximum iterations $K$
[…]
3:for $k=0,1,\dots,K$ do
4:  Train diffusion model $\varepsilon_{\theta}^{(k)}$ on $\mathcal{D}_{\text{train}}^{(k)}$
5:  for all $t\in\mathcal{T}_{\text{target}}\setminus\mathcal{T}_{\text{solved}}^{(k)}$ do
6:   Generate synthetic data $\mathcal{D}_{\text{syn}}^{(k)}(t)$ using $\varepsilon_{\theta}^{(k)}$
7:   Train and evaluate TD3-BC policy $\pi_{\phi}^{(k)}(t)$ on $\mathcal{D}_{\text{syn}}^{(k)}(t)$ to obtain success rate $\mathrm{sr}^{(k)}(t)$
8:  end for
9:  $\mathcal{T}_{\text{solved}}^{(k+1)}\leftarrow\mathcal{T}_{\text{solved}}^{(k)}\cup\{t:\mathrm{sr}^{(k)}(t)>\tau\}$
10:  if $\mathcal{T}_{\text{solved}}^{(k+1)}=\mathcal{T}_{\text{solved}}^{(k)}$ then $c\leftarrow c+1$ else $c\leftarrow 0$ end if
11:  if $c\geq C$ then $\tau\leftarrow\max(\tau-\Delta_{\tau},\tau_{\min})$, $c\leftarrow 0$ end if
12:  $\mathcal{D}_{\text{train}}^{(k+1)}\leftarrow\mathcal{D}\cup\bigcup_{t\in\mathcal{T}_{\text{solved}}^{(k+1)}}\mathcal{D}_{\text{syn}}^{(k)}(t)$
13:  if $\mathcal{T}_{\text{target}}\subseteq\mathcal{T}_{\text{solved}}^{(k+1)}$ then break end if
14:end for
15:return $\varepsilon_{\theta}^{(k)},\{\pi_{\phi}^{(k)}(t)\}_{t\in\mathcal{T}_{\text{target}}},\mathcal{T}_{\text{solved}}^{(k)},\mathcal{D}_{\text{train}}^{(k)}$
4 Experimental Evaluation of the Semantic Compositional Transformer
This section empirically evaluates our compositional transformer for generating robotic data of unseen tasks. Because our algorithm generates millions of synthetic transition tuples for each task, we restrict our experiments to a subset of $64$ tasks of the CompoSuite benchmark, chosen by using one fixed robot arm: IIWA. We consider a setting where we only have training data for $14$ of the $64$ tasks, which we show empirically to be insufficient for zero-shot generalization of a non-compositional data-generating baseline (Appendix A). We use our method to iteratively generate data for the remaining $50$ tasks, and report performance on a test set consisting of $32$ of the $50$ held-out tasks. Note that Algorithm 1 requires online evaluation on the zero-shot tasks, for which we perform 10 trajectory rollouts per task (500 transitions per rollout) every round.
4.1 Baselines
We first compare against two static offline RL approaches, which cannot generate data for new tasks, to demonstrate the value of iterative data generation.
  • Hardcoded Compositional RL   We train the multi-task compositional architecture of Mendez2022 via offline RL using TD3-BC. This architecture was designed specifically to solve CompoSuite tasks, but it employs a hard-coded compositional structure.
  • Semantic Compositional RL   To test the benefits of learned connections in compositional representations, we also implement a semantic compositional RL approach based on our architecture. We use TD3-BC to train a multi-task model which uses our semantic compositional transformer for the encoder. However, rather than decoding each element with its own decoder, we use mean pooling over all output tokens and process the concatenated vector with an additional feed-forward layer to obtain an action.
We then consider three baseline architectures that iteratively generate data per Algorithm 1.
  • Monolithic   To highlight the difficulty of compositional generalization for monolithic architectures, we consider a variant of Synthetic Experience Replay (SynthER; lu2023synthetic). SynthER trains a diffusion model on the data collected by an off-policy RL algorithm (e.g., TD3) to augment the RL batch with artificial transitions. We are specifically interested in the neural network architecture for diffusion, as it has shown promise for generating useful transitions for RL training. In particular, SynthER employs a monolithic architecture that parametrizes the diffusion denoiser $\epsilon_{\theta}$ via several residual feed-forward layers.
    We adapt this architecture to the multi-task setting by conditioning the denoiser $\epsilon_{\theta}$ on the task indicator. At each layer, the noisy transition is augmented by additive embeddings encoding both timestep and task information $\widetilde{x}_{t}=x_{t}+E_{t}(t)+E_{c}(c)$, where $E_{t}(t)$ encodes the diffusion timestep through sinusoidal features and $E_{c}(c)$ linearly projects the multi-hot task indicator into the same latent space. This conditioning strategy injects task information into the residual computation without modifying the architecture.
  • Standard DiT   We then compare against a standard DiT without semantic or compositional tokenization (peebles2023scalable). This DiT simply chops the input into patches of roughly size $15$ and computes the tokens using a shared encoder. This yields a transformer with the same amount of tokens as our compositional semantic encoder but without compositional structure.
  • Semantic DiT   Our tokenization scheme splits the input into semantic patches (e.g., robot state, object state, action) and uses a separate encoder-decoder pair for each element (e.g., one for the IIWA and one for the Jaco). To verify the need to train separate encoder-decoders to learn the different representation spaces for each element, we compare against a DiT that splits the input into semantic patches but trains one shared encoder-decoder across elements of an axis (e.g., one for all robots). While this carries the semantic meaning of the input, it does not model the nodes that constitute the CompoSuite task graph.
In each round of data generation, the diffusion model generates data for the held-out tasks that have not surpassed the threshold $\tau$ (i.e., unseen tasks for the diffusion model). We use the generated data to train task-specific RL policies using TD3-BC for 50,00050,\!000 steps, rolling out $10$ evaluation trajectories every 5,0005,\!000 steps. We keep the best-performing policy for a task across evaluation steps and data generation iterations.
4.2 Zero-shot Generalization
Figure 5: Zero-shot success rate for non-iterable RL models and RL models trained on synthetic data of a generative model at iteration iteration 0, and best zero-shot success rate for RL models trained on synthetic of a generative model after 4 iterations of self-improvement. Overall, the semantic compositional architecture leads to large improvements in the RL as well as generative model setting. (Left) The semantic compositional architecture improves over the hard-coded compositional RL architecture by almost doubling the success rate. (Middle) In the generative model setting, the semantic tokenization approaches beat the architectures that are not adapted to the task. (Right) After 4 rounds of self-improvement, the semantic compositional DiT method achieves highest performance across all other approaches.
To verify the zero-shot abilities of our approach, we pre-train all models on the training tasks and report success on the held-out test tasks. RL baselines directly use a zero-shot policy, while diffusion approaches generate synthetic data and train policies on the generated data. We also run four iterations of iterative self-improvement (Algorithm 1) on diffusion approaches. Performance of RL algorithms is averaged across tasks over $15$ seeds. For generative models, we keep the best-performing policy across the four iterations. We train the diffusion model over three independent seeds, and for each seed we train the policies over five RL training seeds; we report the average across tasks, diffusion seeds, and RL seeds. Error bars indicate standard error across $15$ RL seeds and three diffusion seeds. We report the results in Figure 5.
Reinforcement learning performance   The compositional RL baseline of Mendez2022, which was specifically designed for these tasks, achieves some zero-shot generalization. However, the composition learned by our compositional transformer succeeds twice as often. This provides evidence that our architecture can extract meaningful compositional structure from data. The improved performance suggests that the graph structure learned by our architecture more effectively connects relevant vertices than the hard-coded architecture of Mendez2022.
Initial generative performance   After the first round of training the first diffusion models, RL based on the data generated by the monolithic architecture performs worse than all compositional variants (RL or diffusion), demonstrating the usefulness of composition for efficient zero-shot data generation. The standard DiT model performs worst across all models, indicating that a proper tokenization of the input space is needed. In consequence, any improvement from our method is not a direct result of the transformer being a stronger representation learning architecture. Our semantic compositional data generation process performs nearly on par with the semantic compositional RL baseline. Critically, as we discuss next, the DiT can then generate data for new tasks and iteratively improve its own performance.
Iterated generative performance   Our iterative self-improvement algorithm increases all architectures’ success rates. The monolithic architecture improves by $17\%$, standard DiT by $15\%$, semantic DiT by $16\%$, and semantic compositional DiT by $22\%$—the semantic compositional architecture achieves the largest absolute improvement. These marked improvements indicate that the nature of compositional data is useful for out-of-distribution generation. As our approach can self-improve, it quickly outperforms the static RL baseline without any additional real training data. Note that we can view the threshold $\tau$ as a soft upper bound on success rate, since we generate data that enables as little as $\tau$ success rate per task, and it is challenging to train policies that outperform this level of data quality. With $\tau$ reaching $0.7$ at iteration four and our semantic compositional DiT achieving a success rate of $55\%$, the gap to this soft upper bound closes.
4.3 Iterative Compositional Data Generation
Figure 6: Performance of different diffusion architectures over iterations of our self-improvement procedure. (Left) Zero-shot best success rate achieved so far. RL agents trained on synthetic data from the semantic compositional diffusion architecture consistently achieve higher success rates than agents trained on data from all other diffusion architectures and quickly surpass the semantic compositional RL baseline. (Middle) Number of tasks solved at least once across iterations: RL trained on synthetic data from the semantic compositional diffusion transformer outperforms both baselines and solves nearly all tasks at least once. (Right) Zero-shot best success rate achieved so far, separated by initial task difficulty. Tasks are partitioned into those that exhibit non-zero success at iteration 0 and those that were entirely unsolved. The semantic compositional architecture particularly improves performance on tasks on which it already obtains initial success in iteration 0. Shaded regions indicate standard error over 33 diffusion seeds.
Next, we investigate each round of the iterative procedure for data generation. In every round of Algorithm 1, we evaluate five runs of TD3-BC for each unsolved task ($\text{sr}<\tau$) to average out the randomness from RL training and track the best success rate so far for each task. Figure 6 reports the average success rate in every iteration and the number of tasks that achieve success at least once (i.e. $\text{sr}>0$).
Success rate over time   Figure 6 (left) reiterates the finding that all architectures consistently improve when artificial data is added. All architectures improve at a similar rate, and so the fact that only semantic architectures eventually outperform the RL version of our architecture is largely due to their significantly higher initial success rate. Our compositional semantic architecture only requires one round of self-improvement to exceed the performance of its RL counterpart. This interplay between initial generative performance and downstream RL performance highlights the importance of studying the two in tandem.
Solved tasks over time   Figure 6 (middle) shows that the semantic compositional approach generates data that yields at least one successful trajectory more consistently than the monolithic approach. In addition, after four iterations of refinement, we solve almost every task at least once. This suggests that our approach could serve as a powerful starting point for fine-tuning new policies, since it can drastically reduce the exploration challenge of online RL. Interestingly, while RL using our architecture achieves a higher zero-shot success rate than our generative approach at iteration 0, this success rate is concentrated on a smaller fraction of tasks. This suggests that the diffusion model generalizes more broadly across tasks, but the initial data quality is insufficient to extract good policies in all parts of the state space.
Given that the semantic compositional model yields the most significant improvement in total success but also has a smaller improvement on tasks solved, one might conclude that our architecture is better at iteratively deriving information from successful tasks by refining marginals. To verify this, we analyze whether iterative improvements appear on tasks that see some success or tasks that are not yet solved. In Figure 6 (right), we show that the monolithic architecture improves roughly at the same rate on both the already successful and unsuccessful tasks. While our semantic compositional model also improves in both regimes, it obtains a much larger jump in performance on tasks that see some initial success at iteration 0. In part, this stems from the fact that there are few tasks left on which no success is achieved initially. Yet, it also provides evidence that our semantic compositional model improves encoder-decoder pairs point-wise using self-generated data.
4.4 Analyzing Compositional Structure
This section studies the compositional structure learned by our architecture. As discussed in Section 2.2, we use the Elucidated Diffusion approach (Karras2022), which parameterizes the diffusion process using continuous noise levels $\sigma$ rather than discrete timesteps, with default noise range $\sigma\in[\sigma_{\min},\sigma_{\max}]$. For our analysis, we evaluate the model’s behavior at a noise level $\sigma_{\text{midpoint}}$ corresponding to the midpoint of the generation schedule, computed per the sampling schedule formula (Karras2022):
$\sigma_{i<N}=\left(\sigma_{\max}^{1/\rho}+\frac{i}{N-1}\left(\sigma_{\min}^{1/\rho}-\sigma_{\max}^{1/\rho}\right)\right)^{\rho}\kern 5.0pt,\quad\sigma_{N}=0\kern 5.0pt,$
where $i$ denotes the step. $\sigma_{\text{midpoint}}$ represents a moderate noise level the model encounters during generation. Throughout this section, we use the DiT trained at iteration 0, using only real data.
Intervention influence
Figure 7: Average intervention influence over all training tasks. The heatmap shows how masking each encoder module (rows) changes the predictions produced by each decoder module (columns). Brighter values indicate a larger influence. The plot shows strong diagonal effects, meaning each factor depends most on its own encoder, but it also reveals notable cross-factor interactions. In particular, the robot encoder has a strong effect on several other decoders, especially the reward decoder.
Figure 8: Average attention weights over all training tasks for a single block diffusion transformer. The heatmap shows how much each state component encoder (rows) attends to every other component (columns) when generating denoised representations. Darker colors indicate stronger attention. The plot shows a clear ordering in which the robot component receives the strongest attention from all queries, followed by the goal, the object, and finally the obstacle.
To analyze the compositional dependencies that our model learns, we compute an influence matrix that measures how inputs to each encoder module affect the outputs of each decoder module. For a given task, we generate random Gaussian noise inputs and compute the outputs at $\sigma_{\text{midpoint}}$. We then systematically intervene on each encoder module by zeroing out its output patches and measure the resulting change in each decoder module’s output. The influence of encoder module $i$ on decoder module $j$ is quantified as the $L_{2}$ norm of the normalized difference between the intervened and nominal decoder outputs. Averaging these normalized differences across many noise samples yields an influence matrix whose entries quantify the causal effect of one module on another while remaining comparable across decoders of different dimensionality. This allows us to measure, for example, which predictions are most affected if the object information is missing. Figure 8 presents the average intervention influence matrix over all training tasks.
As expected, the largest deviations happen on the diagonal, as masking a certain element at the encoder makes it difficult to accurately generate that element itself. For example, if the object embedding is masked, the transformer relies exclusively on task conditioning to generate object information. Variations across state components further expose a particular dependency structure among elements. For instance, task input influences object prediction more than it does obstacle prediction. Yet all components depend on each other to some degree. This is not particularly surprising, since the state representation contains relative information (e.g., relative poses). More interestingly, many decoder outputs rely heavily on the robot arm encoding. This highlights the crucial importance of the robot arm in our model for generating data. The largest influence outside the diagonal is for the robot arm input and reward prediction. In general, the reward predictions greatly depend on the state components but not so much on the action, which is correctly inferred by the model since the reward in CompoSuite is only a function of the state. The terminal signal depends exclusively on its own input, since the dataset released by hussing2024robotic contains expert trajectories for almost all tasks, making failure terminations rare in the training data.
Attention masks
Now, we shift our focus to the state encoder structure, the central piece of our architecture design. We study the DiT attention structure by capturing the full 11×11 self-attention matrices. To inspect which encoder outputs map to which decoders, we trained a single-layer transformer. We draw 100 Gaussian inputs, run the model at $\sigma_{\text{midpoint}}$, and compute per-head attention weights from every Multi-Head Self-Attention block. We do this for the 14 training tasks and average across samples and attention heads. We are mostly interested in the state decomposition, which is the main distinguishing feature of our architecture. Figure 8 shows the entries of the attention matrix corresponding to state elements.
The state attention mask reveals that there exists a non-trivial mapping between the state encoder and decoder pairs. First, every decoder pays some attention to its corresponding encoder (the diagonal). Then, we observe an ordering of importance across state elements. Every encoder pays greatest attention to the robot, then the objective, then the object, then the obstacle. The ordering we find is contrary to the hard-coded architecture of Mendez2022, where the robot modules are stacked onto the remaining modules last, implying that other encoders cannot access robot information. This difference may be stem from fundamentally different computations required to learn an RL policy compared to generating RL data.
5 Related Work
Compositional Generalization in Robotics and RL   In robotics, compositional generalization has been pursued through a variety of mechanisms. Some approaches introduce modularity or architectural biases aimed at composing semantic units such as instructions or high-level skills (xu2018ntp; devin2019planarithmetic; kuo2020deepcompositional; wang2023programmatically). Other work targets the control layer directly, designing modular, factorized, or entity-centric policy architectures that encourage reuse of behavioral components across tasks (devin2017learning; mendez2022modular; zhou2025policy). Some approaches seek to automatically identify and decompose policies into functional modules (yang2020multitask; mittal2020learning; goyal2021recurrent). A complementary direction exploits scene-centric formulations that use structured object-relational representations to compose low-level visuomotor skills in novel physical configurations (qi2025composebyfocus).
These approaches demonstrate the importance of leveraging task structure, but they typically assume a hand-designed decomposition of robots, objects, and goals. In contrast, our method learns this structure directly from data by interpreting a diffusion transformer as a GNN and equipping it with factor-specific tokenizers. Whereas prior work uses compositionality to structure policies, we instead use it to structure a generative model of transitions, enabling zero-shot synthesis of data for unseen task compositions.
Generated Data in Robotics and RL    Synthetic data has become a key idea for scaling robotic learning. One line of work expands imitation datasets through trajectory-level augmentations, where expert demonstrations are perturbed, resampled, or regenerated to increase coverage (Mandlekar2023; RoboGen2023; RoCoDA2024; Jiang2025DexMimicGen). Although these methods enrich demonstration sets, they remain constrained to variations of the same underlying tasks without expanding into new compositions.
Complementary efforts in reinforcement learning explore generative replay, where learned generative models synthesize transitions to supplement or replace entries in an agent’s replay buffer (huang2017enhanced; Ludjen2021; Imre2021; lu2023synthetic; voelcker2025madtd). As generative modeling techniques have advanced—from variational autoencoders (kingma2013auto) and generative adversarial networks (goodfellow2014generative) to, more recently, diffusion models (Karras2022)—the fidelity of replayed experience and the sample efficiency of these approaches have improved accordingly. However, these methods still generate data only for the same tasks observed during training. They do not attempt to produce transitions for unseen combinations of factors that fall outside the original task distribution.
Another orthogonal direction emphasizes visual augmentation, including render-driven and vision-only pipelines that procedurally generate synthetic video datasets (Bonetto2023; Singh2024; Yu2024VisualParkour; Re3Sim2024; Real2Render2024), as well as generative and diffusion-based methods that augment images while holding actions constant (chen2023genaug; yu2023scaling). While effective for increasing visual diversity, these approaches do not provide transition-level data reflecting novel task semantics.
Our work is complementary to all of these efforts but differs fundamentally: we generate full state-action-next-state transitions for unseen tasks. Moreover, our iterative procedure evaluates the usefulness of generated data via offline RL, creating a closed-loop mechanism for self-improving compositional data generation.
Compositional Data Generation in Robotics   Recent work has shown that exposing robots to compositional factors of variation can significantly improve generalization and reduce data requirements in manually collected datasets (@gao2024). In parallel, compositional generative models have emerged that synthesize novel object and task combinations to expand the space of training experience (zhou2024robodreamer; barcellona2025dream). These approaches demonstrate the utility of factoring environments into reusable components, but they operate on image representations and often rely on predefined decompositions. More importantly, prior compositional generative approaches do not address the challenge of improving a generative model using its own compositional synthesized data. In contrast, our approach introduces an iterative self-refinement procedure in which the generative model synthesizes transitions for unseen task compositions.
6 Conclusion
In this work, we introduce an iterative compositional data generation framework that uses a semantic compositional diffusion transformer and a self-improvement loop to synthesize and curate manipulation data. This data is of sufficient quality to train policies that solve novel combinations of compositional tasks. Our work shows that compositional data generation can turn limited real interaction into policies that generalize across many tasks. This has the potential to reduce data collection and engineering costs for real-world robotic systems, making it easier to deploy flexible manipulation skills in diverse environments.
At present, our method decides whether a generated task dataset is added to the training pool by running an online evaluation loop. We deploy an RL agent on the newly generated task and include the corresponding data only if its success rate exceeds a fixed threshold. This makes our procedure dependent on online interaction with the environment, which can be costly in many real-world settings. Note that our approach is not tied to this particular choice and any suitable scoring function that assesses the utility of generated data could be used instead. An important direction for future work is to replace this online evaluation with interaction-free or partially offline proxies that can reliably predict the utility of newly generated data.
Acknowledgments
DSB acknowledges support from the Center for Curiosity. MH and EE were partially supported by the Army Research Office under MURI award W911NF201- 0080 and DARPA award HR00112420305. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the view of DARPA, the US Army, or the US government.
Appendix A Regime to Study Compositionality in CompoSuite
Figure 9: Return difference between RL policies trained on data generated by the monolithic architecture and policies trained on ground-truth data over varying number of training tasks. As the number of training tasks approaches 56 $(\sim 20\%)$, there is steep increase in the performance gap indicating the sub-optimality of generated data from the diffusion model.
When sufficient expert data is available, standard feed-forward policies trained with behavioral cloning on the CompoSuite datasets achieve non-trivial zero-shot generalization (hussing2024robotic). However, this assumes access to expert trajectories for hundreds of tasks, which is unrealistic in many robotics applications. As data becomes sparser, exploiting the compositional structure of the tasks becomes more relevant. Here, we verify that the data regime of $14$ training tasks from Sections 4.24.4 is appropriate for studying compositionality. Using all $10$ task-lists from the experimental setup suggested by hussing2024robotic, we construct subsets of training tasks using the first $N$ tasks from each list, where $N\in\{56,98,140,182,224\}$, keeping the set of $32$ test tasks fixed across values of $N$. We then train the SynthER-based architecture introduced in section 4.1 on each subset of training tasks. We generate one million transitions for each test task and train a per-task TD3-BC agent on the generated data. We measure the difference in accumulated return over a set of evaluation trajectories relative to a TD3-BC agent trained on real data. We expect that when the amount of available training data becomes small, the TD3-BC performance should decrease as the generated data quality on out-of-distribution tasks decreases. We report the results in Figure 9.
The results show that when more than $182$ tasks are available for training the diffusion models, the mean gap to the ground-truth policy performance is less than $15\%$. While even the diffusion model trained on $98$ tasks achieves high zero-shot generalization, we see a downward trend below this point. As expected, when we move to $56$ tasks (roughly $20\%$ of the tasks) the performance gap increases drastically, and the model is unable to zero-shot generalize meaningfully. This is a similar data regime to the $14/64$ tasks we used to show the ability of our compositional DiT to learn the underlying graph compositional structure.
Appendix B Additional Experimental Details
This section provides details of the experimental setting used to obtain all results in Section 4 in the main paper.
Computational Requirements
All experiments were conducted on a SLURM-managed GPU cluster equipped with NVIDIA RTX 2080 Ti, RTX 3090, RTX A10, RTX A40, RTX A6000, and L40 GPUs. Training jobs were distributed across these node types, and all model and batch-size configurations were selected to run reliably within the memory constraints of this mixed hardware environment.
B.1 Compositional Data Generation
B.1.1 Diffusion Model Training
We train diffusion models on transition tuples $(s,a,r,s^{\prime},d)$ with total dimension $164$: a 77-dimensional state, an 8-dimensional action, a scalar reward, a 77-dimensional next state, and a binary terminal indicator. All continuous dimensions are standardized (zero mean, unit variance) using statistics computed from the pooled dataset of all training tasks. The terminal indicator remains unnormalized and is discretized with a threshold of $0.5$.
All model variants use the same Elucidated Diffusion schedule; differences arise only from the denoiser architecture and a small number of optimization hyperparameters.
During training, noise levels are sampled from a log-normal distribution with mean $P_{\text{mean}}=-1.2$ and standard deviation $P_{\text{std}}=1.2$. The loss weighting function uses $\sigma_{\text{data}}=1.0$, and the noise schedule spans $\sigma\in[0.002,80]$ with curvature parameter $\rho=7$. Tasks are encoded as 16-dimensional binary indicator vectors, as illustrated in Figure 3, and these task indicators condition the denoiser during both training and generation. The complete set of architectural and optimization hyperparameters for the Monolithic baseline (lu2023synthetic) and our Semantic + Compositional DiT (S+C DiT) model is listed in Table 1.
Table 1: Diffusion model hyperparameters.
COMPONENT MONOLITHIC (lu2023synthetic) S+C DiT (Ours)
Architecture
6-layer MLP,
width 2048
8-layer DiT, hidden size 416, 8 heads
(patch size 15 in standard DiT ablation)
Network capacity 25.88M parameters 26.22M parameters
Batch size 1024 1024
Learning rate $3\times 10^{-4}$ $1\times 10^{-4}$
Weight decay 0.0 0.01
Optimizer AdamW AdamW
LR scheduler Cosine Cosine
Training steps 100,000 100,000
B.1.2 Data Generation
After each diffusion model is trained, we generate synthetic transition datasets for individual tasks using the EMA (Exponential Moving Average) version of the model. Each task is specified as a 4-tuple $(\text{Robot},\text{Object},\text{Obstacle},\text{Objective})$ and encoded using the same 16-dimensional task indicator from Figure 3. The Monolithic and S+C DiT pipelines use identical sampling configurations, except for the generator batch size, which is reduced for S+C DiT to satisfy GPU memory constraints. The hyperparameters used for synthetic data generation are summarized in Table 2.
Table 2: Synthetic data generation hyperparameters.
COMPONENT MONOLITHIC (lu2023synthetic) S+C DiT (Ours)
Generated samples per task 1,000,000 1,000,000
Sampling steps 128 128
Noise perturbation strength 80 80
Minimum noise level for perturbation 0.05 0.05
Maximum noise level for perturbation 50 50
Relative perturbation noise scale 1.003 1.003
Generator batch size 100,000 25,000
Batches per task 10 40
B.1.3 Policy Training
We train TD3-BC policies (Fujimoto2021Minimalist) on the synthetic transition datasets generated for each task. The same TD3-BC configuration is used across all experiments, including both the monolithic and S+C DiT pipelines and all iterations. The only exception is the compositional RL baselines, which use different compositional policy architectures that are described in Appendix B.2. All policies are trained offline on synthetic data and are then evaluated online in the corresponding CompoSuite environment.
For all test tasks, we train policies using five random seeds and report the mean success rate. States are normalized using the mean and standard deviation computed from each task’s synthetic training dataset. The complete set of TD3-BC hyperparameters used in all experiments is provided in Table 3.
Table 3: TD3-BC offline RL training hyperparameters.
COMPONENT VALUE
Algorithm TD3-BC
Actor network MLP with 2 hidden layers of width 256
Critic networks MLP with 2 hidden layers of width 256
Learning rate (actor, critics) $3\times 10^{-4}$
Optimizer Adam
Batch size 1024
Training steps 50,000
Discount factor ($\gamma$) 0.99
Target network update ($\tau$) 0.005
Regularization coefficient ($\alpha$) 2.5
Policy noise 0.2
Noise clip 0.5
Policy update frequency 2
Evaluation frequency 5,000 steps
Evaluation episodes 10
State normalization Yes
Reward normalization No
Training seeds 044
B.1.4 Iterative Bootstrapping
The iterative bootstrapping procedure follows Algorithm 1. We initialize the success threshold at $\tau_{0}=0.8$. If no new tasks satisfy this threshold for one iteration ($C=1$), the threshold is automatically reduced by $\Delta_{\tau}=0.1$, with a lower bound of $\tau_{\text{min}}=0.5$. The IIWA-only task list used for this experiment is described in Appendix B.3. The complete set of hyperparameters for this procedure is reported in Table 4.
Table 4: Compositional iterative bootstrapping hyperparameters.
COMPONENT VALUE
Initial success threshold ($\tau_{0}$) 0.8
Minimum threshold ($\tau_{\text{min}}$) 0.5
Threshold reduction amount ($\Delta_{\tau}$) 0.1
Patience ($C$) 1 iteration
Training tasks ($|\mathcal{T}\text{train}|$) 56 of 256 tasks, or 14 of 64 IIWA-only tasks
Diffusion seeds 022
B.2 Compositional RL Baselines
We compare against two compositional RL baselines that train multitask TD3-BC policies on expert demonstrations from the same 14 training tasks used for diffusion training. Their train/test split matches the diffusion setup to allow a fair zero-shot generalization comparison. The list of training and test tasks is described in Appendix B.3. All baselines are evaluated on the corresponding 32 held-out test tasks using $15$ random seeds. Both baselines follow the TD3-BC configuration in Table 3, but differ in three ways: (1) they train on a multitask dataset that combines demonstrations from all 14 training tasks, (2) they employ compositional policy architectures, and (3) their batch sizes are scaled with the number of training tasks (i.e., a multiple of the number of tasks), following the strategy used in Mendez2022.
Table 5: Compositional RL baseline hyperparameters for multitask TD3-BC training.
COMPONENT HC RL (Mendez2022) S+C RL (Ours)
Algorithm TD3-BC TD3-BC
Training tasks 14 14
Test tasks 32 32
State dimension 93 (with task IDs) 93 (with task IDs)
Action dimension 8 8
Actor architecture
Hardcoded compositional
MLP
Semantic compositional
transformer
Actor output dimension 8 8
Critic architecture
Hardcoded compositional
MLP (state+action)
Semantic compositional
transformer (state+action)
Critic output dimension 1 1
Compositional module sizes $(32),(32,32),(64,64,64),(64,64,64)$
Compositional hierarchy Obstacle $\rightarrow$ Object $\rightarrow$ Subtask $\rightarrow$ Robot
Transformer hidden size 72
Transformer depth 1
Transformer heads 4
Transformer MLP ratio 1.20
Policy network capacity 107.94K parameters 106.29K parameters
Learning rate $3\times 10^{-4}$ $3\times 10^{-4}$
Optimizer Adam Adam
Batch size 3584 (14 tasks × 256) 1792 (14 tasks × 128)
Training steps 50,000 50,000
Discount factor ($\gamma$) 0.99 0.99
Target network update ($\tau$) 0.005 0.005
Regularization coefficient ($\alpha$) 2.5 2.5
Policy noise 0.2 0.2
Noise clip 0.5 0.5
Policy update frequency 2 2
Evaluation frequency 5,000 steps 5,000 steps
Evaluation episodes 10 per test task 10 per test task
State normalization Yes Yes
Reward normalization No No
Training seeds 01414 01414
The Hardcoded Compositional RL (HC RL) baseline uses the modular architecture of Mendez2022 with component-specific networks for each task element. The hardcoded architecture follows a hierarchical graph structure with the ordering Obstacle $\rightarrow$ Object $\rightarrow$ Subtask $\rightarrow$ Robot. Hidden layer sizes are $(32)$ for Obstacle, $(32,32)$ for Object, $(64,64,64)$ for Subtask, and $(64,64,64)$ for Robot.
The Semantic Compositional RL (S+C RL) baseline uses a transformer with semantic tokens corresponding to the object, obstacle, goal, and robot. Compositional encoders produce token embeddings, and task conditioning is applied using Adaptive Layer Normalization (AdaLN). The transformer uses hidden size 72, depth 1, 4 attention heads, MLP ratio 1.20, and no dropout. For this baseline, the batch size is further reduced by one half compared to Mendez2022 to satisfy GPU memory constraints.
The full set of hyperparameters is listed in Table 5.
B.3 Task List
For the results in Appendix A, we use the ten task lists released by hussing2024robotic. For the IIWA-only experiments in Section 4, we construct a train/test split over the full IIWA task space, defined by all combinations of the IIWA robot with:
  • Object: Box, Dumbbell, Hollowbox, Plate,
  • Obstacle: GoalWall, None, ObjectDoor, ObjectWall,
  • Objective: PickPlace, Push, Shelf, Trashcan.
This yields $4\times 4\times 4=64$ tasks. We generate a random split using seed 0, producing 32 training and 32 test tasks with no overlap. We use the first 14 training tasks for all diffusion and multitask policy experiments, and evaluate on all 32 held-out test tasks. Table 6 lists the tasks, and Figures 10 and 11 visualize the split.
Figure 10: Visualization of the 14 training tasks used in the IIWA-only split. Tasks are shown in numerical order (1–14), arranged left-to-right and top-to-bottom. Each image depicts one unique combination of Object, Obstacle, and Objective paired with the IIWA robot.
Figure 11: Visualization of the 32 held-out test tasks used for zero-shot evaluation. Tasks are displayed in numerical order (1–32), arranged left-to-right and top-to-bottom. Each image corresponds to a distinct unseen combination of Object, Obstacle, and Objective in the IIWA environment.
Table 6: Training and test tasks used for IIWA-only experiments.
Training Tasks (14) Test Tasks (32)
1. IIWA, Box, ObjectDoor, Trashcan
2. IIWA, Hollowbox, ObjectDoor, PickPlace
3. IIWA, Dumbbell, ObjectDoor, PickPlace
4. IIWA, Dumbbell, ObjectWall, Push
5. IIWA, Plate, None, Shelf
6. IIWA, Box, GoalWall, Trashcan
7. IIWA, Plate, ObjectWall, Shelf
8. IIWA, Hollowbox, GoalWall, Trashcan
9. IIWA, Box, ObjectWall, Shelf
10. IIWA, Box, None, Trashcan
11. IIWA, Plate, ObjectWall, PickPlace
12. IIWA, Box, GoalWall, PickPlace
13. IIWA, Box, None, Push
14. IIWA, Box, ObjectDoor, Shelf
1. IIWA, Dumbbell, GoalWall, Shelf
2. IIWA, Box, None, PickPlace
3. IIWA, Box, GoalWall, Shelf
4. IIWA, Hollowbox, None, PickPlace
5. IIWA, Dumbbell, ObjectDoor, Push
6. IIWA, Box, None, Shelf
7. IIWA, Plate, None, PickPlace
8. IIWA, Dumbbell, None, Shelf
9. IIWA, Dumbbell, ObjectDoor, Shelf
10. IIWA, Hollowbox, GoalWall, PickPlace
11. IIWA, Dumbbell, GoalWall, Trashcan
12. IIWA, Plate, ObjectDoor, Push
13. IIWA, Plate, ObjectDoor, Shelf
14. IIWA, Hollowbox, None, Trashcan
15. IIWA, Box, ObjectDoor, PickPlace
16. IIWA, Box, ObjectDoor, Push
17. IIWA, Hollowbox, None, Shelf
18. IIWA, Dumbbell, ObjectWall, Shelf
19. IIWA, Hollowbox, GoalWall, Shelf
20. IIWA, Box, ObjectWall, Push
21. IIWA, Hollowbox, ObjectWall, Shelf
22. IIWA, Hollowbox, None, Push
23. IIWA, Plate, GoalWall, Shelf
24. IIWA, Plate, ObjectDoor, PickPlace
25. IIWA, Plate, GoalWall, Trashcan
26. IIWA, Dumbbell, GoalWall, PickPlace
27. IIWA, Hollowbox, ObjectDoor, Trashcan
28. IIWA, Dumbbell, ObjectWall, Trashcan
29. IIWA, Plate, None, Push
30. IIWA, Plate, GoalWall, Push
31. IIWA, Dumbbell, None, Push
32. IIWA, Plate, GoalWall, PickPlace